The monoatomic ideal gas receives 750 J of heat energy (heat) while working on an environment of 625 J. There are 1.3 moles of gas in the sample. Find the difference between the final and initial temperature of the system. Appropriate constants taken from tables. (Answer: ΔT = 7.71 K)


1
Expert's answer
2021-04-16T07:23:29-0400

According to the first law of thermodynamics,


"Q=\Delta U+W,\\\space\\
\Delta U=Q-W=\frac32 nR\Delta T,\\\space\\
\Delta T=\frac{2(Q-W)}{3nR}=\frac{2(750-625)}{3·1.3·8.314}=7.71\text{ K}."